OpenStudy (anonymous):

second to last: use Cramer's Rule to solve the system of equations:

OpenStudy (anonymous):

Cramer's Rule: \[-A ^{-1}D=X\]

OpenStudy (anonymous):

system: x-y+2z+1=0 3x+y+4z-2=0 2y+z-2=0

OpenStudy (anonymous):

so we need A, D, detA and A^-1

OpenStudy (anonymous):

how do you draw matrices?

OpenStudy (anonymous):

using the equation thing?

OpenStudy (anonymous):